At a certain temperature, the K_p for the decompos
Answer #1

So we have got the partial pressure of H2 and S at equilibrium

Partial Pressure of H2 = S = 0.121 atm

Partial Pressure of H2S at equilibrium = 0.141 - 0.121 = 0.02 atm

So total pressure at equilbrium = 2 * 0.121 + 0.02 = 0.242 + 0.02 = 0.262 atm

Thank You